OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 14 x + 3y = 9 (1, 12) (−1, −12) (12, −1) (12, 1)

OpenStudy (anonymous):

(12,-1)

hero (hero):

x = 14 + 2y x = 9 - 3y x = x 14 + 2y = 9 - 3y 3y + 2y = 9 - 14 5y = -5 y = -1 x = 12

hero (hero):

therefore, (12,-1)